# Can Large Language Models Identify And Reason About Security Vulnerabilities? Not Yet

### v1

@@ 3.6 Evaluator
The output of an LLM for a specific test is passed to the Evaluator (see Figure 2). As our framework is fully automated, we leverage GPT-4 to analyze the response. First, the response is passed to GPT-4 333Our manual evaluation shows that GPT-4 performs the best for this extraction process., with an additional role-based instruction prompt (shown in Figure 5 in the Appendix) in the ‘system’ input to extract two pieces of information from the raw response. The first one is the binary answer, which is “yes/no” based on whether the LLM found a vulnerability in the given code or not. We find that in some cases the LLM does not provide a definite answer, therefore we include a third verdict, i.e., “n/a.”444For example, in some cases LLMs provide some variation of the following response: “As an AI model I cannot answer this question.” The second part of the information extracted is the textual reasoning provided by the LLM (). To extract it, we ask GPT-4 to summarize the reason described by the model output on why a vulnerability is present in the code or not, in 100 words. We extract the summary of root cause of vulnerability from raw responses of LLMs to maintain consistency for further evaluation methods, and to avoid contents like suggestions, fixed code, etc. to be analyzed. In the rest of this section we provide more details on the evaluation metrics used to evaluate the final answer, and its reasoning provided by LLM for the vulnerability detection task.
@@
Cosine Similarity is a metric commonly used in NLP to measure how similar two documents are irrespective of their sizes. We first convert the summaries and into fixed length vectors using OpenAI’s embedding model ‘text-similarity-davinci-001’ and calculate the cosine similarity between them. We find that the optimal threshold is , and consider two summaries similar if their cosine similarity exceeds this value.
@@ 4.1 Evaluation for Deterministic Responses
| chat-bison |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
2/10 | 8/10 |
10/10 |
0/10 |
| codechat-bison | 9/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
0/10 |
9/10 |
0/10 |
10/10 |
0/10 |
10/10 |
| codellama34b |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
5/10 | 7/10 |
| gpt-3.5 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
10/10 |
0/10 |
| gpt-4 |
0/10 |
10/10 |
6/10 |
10/10 |
1/10 |
10/10 |
6/10 |
10/10 |
0/10 |
10/10 |
7/10 |
10/10 |
@@
| chat-bison |
10/10 |
10/10 |
10/10 |
8/10 |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
10/10 |
0/10 |
10/10 |
| codechat-bison |
10/10 |
10/10 |
10/10 |
2/10 |
10/10 |
7/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
10/10 |
| codellama34b |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
0/10 |
| gpt-3.5 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
1/10 |
10/10 |
10/10 |
10/10 |
7/10 |
| gpt-4 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
5/10 |
@@
| chat-bison |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
0/10 |
10/10 |
9/10 |
0/10 |
| codechat-bison |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
| codellama34b |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
0/10 |
10/10 |
| gpt-3.5 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
10/10 |
0/10 |
| gpt-4 |
0/10 |
10/10 |
8/10 |
10/10 |
2/10 |
10/10 |
4/10 | 9/10 |
0/10 |
10/10 |
4/10 |
10/10 |
@@
| chat-bison |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
10/10 |
| codechat-bison |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
0/10 |
| codellama34b |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
0/10 |
| gpt-3.5 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
10/10 |
| gpt-4 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
0/10 |
10/10 |
10/10 |
10/10 |
10/10 |
10/10 |
9/10 |
@@ 4.2 Performance Over Range of Parameters
As running experiments on all prompts is not feasible due to budget constraints, we perform the experiment on only one prompt, S4 (ZS-RO), based on three reasons; (1) this prompt does not provide any additional information like definition or step-by-step reasoning instructions to the LLM, so the response will be mainly based on the intrinsic knowledge of the model from its training data and no external instruction or information will influence its reasoning or creativity at a higher temperature, (2) being a zero-shot prompt it does not limit the creativity or reasoning of the model as in few-shot prompting might influence the model to mimic the few-shot examples, and (3) this prompt is the most non-deterministic zero-shot prompt as shown in Tables VIII(b) and VII(b), allowing it to show more randomness or creativity at higher temperature. We evaluate LLMs on six temperature values: their recommended value , and over a range of values from 0 to 1 i.e., , , , , . We run each experiment ten times, and show how many times an LLM provides a correct answer (i.e., accuracy) and correct reasoning (i.e., reasoning score). The results are summarized in Tables IX and X.
@@ 4.4 Faithful Reasoning
Faithful reasoning is the quality of an LLM to provide the right reasoning for the right answer or vice versa [51]. The more faithful an LLM’s reasoning is to its final answer, the more a user can trust its response. Table XI shows that even when they provide the right response, LLMs sometimes provide the wrong reason for this decision. In this section, we further analyze the faithful reasoning of LLMs on their decisions from the experiment discussed in the previous section, focusing on four aspects: (1) for how many cases does the LLM provide a reasoning for the presence of a vulnerability in a code snippet at all, (2) for how many correct answers the LLM also provides a correct reasoning, (3) for how many correct answers the LLM provides an incorrect reasoning, and (4) for how many answers does the LLM provide the wrong answer, but a correct reasoning. Table XII provides an overview of the results of this experiment.
Observations. While in the vast majority of cases the answer and reasoning for the tested LLMs align, every LLM presents cases where it provides the correct reasoning but this leads to a wrong answer (as shown in Figure 3), similarly, we find cases where an LLM provides the right answer but its reasoning or root cause is not correct (as shown under ‘codechat-bison (1st Response)’ in Figure LABEL:fig:output-inconsistency). We also find that Google’s PaLM2 models have overall lower reasoning rate as they do not explain their decisions in many cases, while GPT models show comparatively higher reasoning rates, especially ‘gpt-4’ and ‘codellama34b’ provide a reason for every answer. Our results cast doubts on the reasoning capabilities of current LLMs, indicating that in many cases their answers may not be backed by thorough and correct reasoning.
@@ 4.5 Evaluation Over Variety of Vulnerabilities
![[Uncaptioned image]](x6.png)
Observations. Most models show poor performance in classifying the patched versions correctly, which makes these LLMs non-suitable for real-world cases as they will mostly flag safe code as vulnerable, causing manyfalse alarms. We observe that few-shot prompting performs better than zero-shot prompting for almost all models, and role-oriented prompts almost always perform better than task-oriented prompts. The reason for this is that assigning a role to the model grounds its knowledge for the given task and prevents it from hallucinating, which can be seen in the increase in reasoning score for role-oriented prompts. However, Table XIII shows that ‘codechat-bison@001’ does not provide answers for ‘CWE-787’ and ‘CWE-416’ for few-shot prompts.
@@ 4.7 Robustness to Code Augmentations
In this section we test the robustness of LLMs by testing them against the code augmentations described in Section V. Our results are summarized in Table VI. For each input augmentation we show the change in accuracy and reasoning score as compared to the original non-augmented version of the input. For each LLM, we test each augmentation using three prompts: standard prompt ‘S1,’ and the best zero-shot (ZS) and few-shot (FS) prompts.666Since we show in Sections 4.5 and 4.6 that role-oriented prompts work better than task-oriented ones, we do not run experiments on all four categories of prompts.
@@ 4.8 Real-World Cases
![[Uncaptioned image]](x11.png)
![[Uncaptioned image]](x12.png)
Observations. Overall, LLMs do not seem to generalize for detecting vulnerabilities in real-world CVEs that are outside their knowledge cut-off date, with all studied LLMs providing incorrect answers for several of our test cases. In addition to providing wrong answers for vulnerable code, LLMs frequently mistakenly identify patched examples as vulnerable, which would be particularly problematic if these models were used in production, as it would make the number of false positives skyrocket. We also observe that few-shot prompting does not work in case of real-world scenarios, likely because LLMs fail to extrapolate information from the provided examples (despite being from the same CWEs) and apply it to other software code-bases. At the same time, we find that the zero-shot role-oriented prompt ‘R2’ shows relatively better performance for all LLMs , which indicates that grounding LLM’s role as ‘security expert’ and providing them explicit guidelines to follow a human-like multi-step vulnerability detection process can improve performance, but is still insufficient for real-world deployment.
@@ 5 Discussion
Our fully automated evaluation framework allows users to evaluate any chat-based LLM for its ability to identify software vulnerabilities. Further, our framework will be publicly released so that the community will be able to evaluate LLMs released in the future. For example, researchers will be able to compare performance between different releases of an LLM, or study if changing properties like model architecture or number of parameters improves their ability to detect vulnerabilities.777We include these case studies for ‘codellama7B’, ‘codellama13B’, and ‘starchat-beta’ in tables in Appendix
@@
Answer and Reason Extraction. We use GPT-4 to parse the LLM output and extract the final answer and reasoning, using a prompt that requires GPT-4 to answer in a given format (as shown in Figure 5 in the Appendix), otherwise further steps of our analysis would fail. We manually analyze 100 extracted answers and reasonings by GPT-4 and only two were not answered in the given format.888We note that using the newest ‘GPT-4 Turbo,’ which provides responses in json format, could eliminate even these two anomalies.
Knowledge Cut-Off. To evaluate newer LLMs, researchers might have to identify CVEs that were release after their knowledge cutoff, to avoid biases in the results. Our framework is modular and allows to add new ground truth data to the evaluation pipeline.
@@ Acknowledgments
We would like to thank Syed Qasim and Pujan Paudel for their help in generating ground-truth reasoning for the code scenarios. This work was supported by the Red Hat Collaboratory and by the NSF under grants CNS-1942610 and CNS-2127232, respectively. Any opinions, findings, and conclusions, or recommendations expressed are those of the authors and do not necessarily reflect the views of the sponsors.

### v2

@@ 3.6 Evaluator
The output of an LLM for a specific test is passed to the Evaluator (see Figure 2). As SecLLMHolmes is fully automated, we leverage GPT-4 to analyze the response. First, the response is passed to GPT-4 444Our manual evaluation shows that GPT-4 performs the best for this extraction process., with an additional role-based instruction prompt (shown in Figure 7 in the Appendix) in the ‘system’ input to extract two pieces of information from the raw response. The first one is the binary answer, which is “yes/no” based on whether the LLM found a vulnerability in the given code or not. We find that in some cases the LLM does not provide a definite answer, therefore we include a third verdict, i.e., “n/a.”555For example, in some cases LLMs provide some variation of the following response: “As an AI model I cannot answer this question.” The second part of the information extracted is the textual reasoning provided by the LLM (). To extract it, we ask GPT-4 to summarize the reason described by the model output on why a vulnerability is present in the code or not, in 100 words. We extract the summary of root cause of vulnerability from raw responses of LLMs to maintain consistency for further evaluation methods, and to avoid contents like suggestions, fixed code, etc. to be analyzed. In the rest of this section we provide more details on the evaluation metrics used to evaluate the final answer, and its reasoning provided by LLM for the vulnerability detection task.
@@
Cosine Similarity is a metric commonly used in NLP to measure how similar two documents are irrespective of their sizes. We first convert the summaries and into fixed length vectors using OpenAI’s embedding model ‘text-similarity-davinci-001’ and calculate the cosine similarity between them. Similar to , we find that the optimal threshold is , and consider two summaries similar if their cosine similarity exceeds this value.
@@ 4.1 Evaluation for Deterministic Responses
| chat-bison | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 2/10 | 8/10 | 10/10 | 0/10 |
| codechat-bison | 9/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 0/10 | 9/10 | 0/10 | 10/10 | 0/10 | 10/10 |
| codellama34b | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 5/10 | 7/10 |
| gpt-3.5 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 10/10 | 0/10 |
| gpt-4 | 0/10 | 10/10 | 6/10 | 10/10 | 1/10 | 10/10 | 6/10 | 10/10 | 0/10 | 10/10 | 7/10 | 10/10 |
@@
| chat-bison | 10/10 | 10/10 | 10/10 | 8/10 | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 10/10 | 0/10 | 10/10 |
| codechat-bison | 10/10 | 10/10 | 10/10 | 2/10 | 10/10 | 7/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 10/10 |
| codellama34b | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 0/10 |
| gpt-3.5 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 1/10 | 10/10 | 10/10 | 10/10 | 7/10 |
| gpt-4 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 5/10 |
@@
| chat-bison | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 0/10 | 10/10 | 9/10 | 0/10 |
| codechat-bison | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 |
| codellama34b | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 0/10 | 10/10 |
| gpt-3.5 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 10/10 | 0/10 |
| gpt-4 | 0/10 | 10/10 | 8/10 | 10/10 | 2/10 | 10/10 | 4/10 | 9/10 | 0/10 | 10/10 | 4/10 | 10/10 |
@@
| chat-bison | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 10/10 |
| codechat-bison | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 0/10 |
| codellama34b | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 0/10 |
| gpt-3.5 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 10/10 |
| gpt-4 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 0/10 | 10/10 | 10/10 | 10/10 | 10/10 | 10/10 | 9/10 |
@@ 4.2 Performance Over Range of Parameters
As running experiments on all prompts is not feasible due to budget constraints, we perform the experiment on only one prompt, S4 (ZS-RO), based on three reasons; (1) this prompt does not provide any additional information like definition or step-by-step reasoning instructions to the LLM, so the response will be mainly based on the intrinsic knowledge of the model from its training data and no external instruction or information will influence its reasoning or creativity at a higher temperature, (2) being a zero-shot prompt it does not limit the creativity or reasoning of the model as in few-shot prompting might influence the model to mimic the few-shot examples, and (3) this prompt is the most non-deterministic zero-shot prompt as shown in Tables VIII(b) and VII(b), allowing it to show more randomness or creativity at higher temperature. We evaluate LLMs on six temperature values: their recommended value , and over a range of values from 0 to 1 i.e., , , , , . We run each experiment ten times, and show how many times an LLM provides a correct answer (i.e., accuracy) and correct reasoning (i.e., reasoning score). The results are summarized in Tables IX(d) and X(d).
@@ 4.4 Faithful Reasoning
Faithful reasoning is the quality of an LLM to provide the right reasoning for the right answer or vice versa [50]. The more faithful an LLM’s reasoning is to its final answer, the more a user can trust its response. Table XI shows that even when they provide the right response, LLMs sometimes provide the wrong reason for this decision. In this section, we further analyze the faithful reasoning of LLMs on their decisions from the experiment discussed in the previous section, focusing on five aspects: (1) for how many cases does the LLM provide a reasoning for the presence of a vulnerability in a code snippet at all, (2) for how many correct answers the LLM also provides a correct reasoning, (3) for how many correct answers the LLM provides an incorrect reasoning, (4) for how many incorrect answers the LLM provides an incorrect and (5) for how many answers does the LLM provide the wrong answer, but a correct reasoning. Table XII provides an overview of the results of this experiment.
Observations. While in the vast majority of cases the answer and reasoning for the tested LLMs align, every LLM presents cases where it provides the correct reasoning but this leads to a wrong answer (as shown in Figure 4). Similarly, we find cases where an LLM provides the right answer but its reasoning or root cause is not correct (as shown under ‘codechat-bison (1st Response)’ in Figure LABEL:fig:output-inconsistency). We also find that Google’s PaLM2 models have overall lower reasoning rate as they do not explain their decisions in many cases, while GPT models show comparatively higher reasoning rates, especially ‘gpt-4’ and ‘codellama34b’ provide a reason for every answer. Our findings suggest that, in certain cases, current LLMs’ responses might not fully rely on faithful and accurate reasoning.
@@ 4.5 Evaluation Over Variety of Vulnerabilities
Observations. Most models show poor performance in classifying the patched versions correctly, which makes these LLMs non-suitable for real-world cases as they will mostly flag safe code as vulnerable, causing manyfalse alarms. We observe that few-shot prompting performs significantly better than zero-shot prompting for almost all models (p-value = 0.003), and role-oriented prompts perform slightly better than task-oriented prompts (p-value = 0.1). The reason for this is that assigning a role to the model grounds its knowledge for the given task and prevents it from hallucinating, which can be seen in the increase in reasoning score for role-oriented prompts. However, Table XIII shows that ‘codechat-bison@001’ does not provide answers for ‘CWE-787’ and ‘CWE-416’ for few-shot prompts.
@@ 4.7 Robustness to Code Augmentations
In this section we test the robustness of LLMs by testing them against the code augmentations described in Section V. Our results are summarized in Table VI. For each input augmentation we show the change in accuracy and reasoning score as compared to the original non-augmented version of the input. For each LLM, we test each augmentation using three prompts: standard prompt ‘S1,’ and the best zero-shot (ZS) and few-shot (FS) prompts.777Since we show in Sections 4.5 and 4.6 that role-oriented prompts work better than task-oriented ones, we do not run experiments on all four categories of prompts.
@@ 4.8 Real-World Cases
Observations. Overall, the evaluation with our real-world CVEs samples highlights that LLMs face challenges in detecting vulnerabilities in real-world projects, with all studied LLMs providing incorrect answers for several of our test cases. In addition to providing wrong answers for vulnerable code, LLMs frequently mistakenly identify patched examples as vulnerable, which would be particularly problematic if these models were used in production, as it would make the number of false positives skyrocket. We also observe that few-shot prompting does not work in case of real-world scenarios, likely because LLMs fail to extrapolate information from the provided examples (despite being from the same CWEs) and apply it to other software code-bases. At the same time, we find that the zero-shot role-oriented prompt ‘R2’ shows relatively better performance for all LLMs , which indicates that grounding LLM’s role as ‘security expert’ and providing them explicit guidelines to follow a human-like multi-step vulnerability detection process can improve performance, but is still insufficient for real-world deployment.
@@ 5 Discussion
SecLLMHolmes allows users to evaluate any chat-based LLM for its ability to identify software vulnerabilities. Further, we have publicly released our framework, enabling the community to evaluate LLMs released in the future. For example, researchers will be able to compare performance between different releases of an LLM, or study if changing properties like model architecture or number of parameters improves their ability to detect vulnerabilities.888We include these case studies for ‘codellama7B’, ‘codellama13B’, and ‘starchat-beta’ in tables in Appendix
@@
Answer and Reason Extraction. We use GPT-4 to parse the LLM output and extract the final answer and reasoning, using a prompt that requires GPT-4 to answer in a given format (as shown in Figure 7 in the Appendix), otherwise further steps of our analysis would fail. We manually analyze 100 extracted answers and reasonings by GPT-4 and only two were not answered in the given format.999We note that using the newest ‘GPT-4 Turbo,’ which provides responses in json format, could eliminate even these two anomalies.
Knowledge Cut-Off. To evaluate newer LLMs, researchers might have to identify CVEs that were release after their knowledge cut-off, to avoid biases in the results. Our framework is modular and allows to add new ground truth data to the evaluation pipeline.
@@ Acknowledgments
We would like to thank Syed Qasim and Pujan Paudel for their help in generating ground-truth reasoning for the code scenarios. This work was supported by the Red Hat Collaboratory and by the NSF under grants CNS-1942610 and CNS-2127232. Any opinions, findings, and conclusions, or recommendations expressed are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Appendix A Examples of Code Difficulty Levels
Appendix B Meta-Review
The following meta-review was prepared by the program committee for the 2024 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
B.1 Summary
This paper suggests an automated framework for evaluating LLMs (Large Language Models) for identifying and reasoning about security vulnerabilities. The authors conducted experiments on various options of LLMs (e.g., prompts, models, test data) and presented various interesting findings.
B.2 Scientific Contributions
Provides a Valuable Step Forward in an Established Field
Creates a New Tool to Enable Future Science
Provides a New Data Set For Public Use
B.3 Reasons for Acceptance
1.
This paper suggests a fully automated framework to evaluate various aspects of LLM’s capability in identifying vulnerabilities, which is gaining more attention from the community.
2.
The paper examines several aspects that have been less studied previously, including code complexity, vulnerability reasoning, and use of chat-based LLMs (e.g., GPT-4) in vulnerability detection.
3.
The authors have developed a new benchmark consisting of 228 code scenarios. This includes 48 hand-crafted examples (starting from most critical CWEs), 30 examples from real-world CVEs (taken from open source projects) and 150 codes obtained by trivial and non-trivial augmentation. If released, this dataset can help in carrying out further research, allowing an alignment for future analysis.
B.4 Noteworthy Concerns
1.
Some findings seem well-known to the community (e.g., non-determinism, and non-robustness by specific code augmentations).
2.
The size of the data set is still limited, especially for the real-world CVEs (the data set only includes 15 CVEs now). Moreover, the dataset includes many augmented codes (150/228). The small data set may lead to significant bias in the measurement results, making it less reliable.
Appendix C Response to the Meta-Review
We thank the reviewers for their valuable feedback and the shepherd for helping us improve the paper. We would like to respond to the meta-review concerns as follows:
1.
While instances of non-determinism and non-robustness to specific code augmentations have been found in previous work, to the best of our knowledge we are the first ones to systematically evaluate these issues in the context of vulnerability detection.
2.
The size of our dataset is comparable to previous work like [8] which only used 12 CVEs. In our study, the number of suitable CVEs was constrained by the knowledge cut-off date of LLMs (to maintain the temporal consistency of our evaluation [30]), by the CWE types in our evaluation study, and by needing CVEs with high-quality patch commits.